OpenStudy (anonymous):

4 = -1/4x + 3 solving for x. HELP :(

OpenStudy (tkhunny):

Have you considered first multiplying by 4?

OpenStudy (anonymous):

is it (-1/4x) or -1/(4x+3)

OpenStudy (anonymous):

4 = -1/4x +3 1=-1/4x -4 = x

OpenStudy (anonymous):

thanks kelkoo